# Terrain-aware Low Altitude Path Planning

### v1

@@ 3 Simulation Experiments
Figure 6: Example of collected images.
In this section, we explain the setup of the simulation experiments and demonstrate the results.
@@ 3.1 Terrain
We build a custom terrain using Blender [24]. The terrain with the corresponding point cloud are visualized in Fig 5, The terrain is exported to NVIDIA Isaac Sim to perform simulation experiments. The size of the terrain is about 820m $\times$ 820m $\times$ 60m, which is scaled down from its original size to accelerate rendering and data collection process.
@@ 3.2 Expert Planner Parameters
We set $\alpha=1.0,\beta=0.1$ in Eq 2.3. The minimal elevation with respect to the terrain is set to $5.0$m and the maximal elevation is set to $15.0$m. Since our terrain is scaled down, we accordingly scale down the minimum turning radius to $3$m and set the maximum climbing rate to $0.6$ rad/s. The cruise speed is set to $5$m/s. The range for RRT* is decreased to $300$m. The resolution used to check state validity is increased to $0.01$.
@@ 3.3 Dataset Preparation
We collect expert demonstrations on 1/4 of the terrain and use the full terrain for testing. We collect 1200 trajectories where the start locations and goal locations are randomly sampled such that each pair of start and goal locations are at least separated by 50m.
@@
To keep the inputs at similar scale, the heading vectors are divided by $100$, the log depth images are divided by $10.0$ and RGB images are divided by $255$.
@@ 3.5 Baseline Policy
In addition, we also train a baseline policy for comparison study, which uses the standard behavior cloning (BC) approach. In other words, the baseline policy does not have the loss terms $L_{\text{terrain}}$ and $L_{\text{elevation}}$. And the loss term $L_{\text{bc}}$ applies to all the three coordinates $x,y,z$ instead of just $x,y$. The other setups are completely the same (i.e. dataset, architecture, and hyperparameters).
@@ 3.6 Results
To compare with the baseline policy, which we term as BC, we select 20 pairs of start and goal locations that are separated by no more than 100m (since the policies were trained on heading vectors with norms less than the path length equivalent to $H$ steps, which is around $112.5$m). The start and goal locations are selected only in the unseen part of the terrain. Since both BC and our policy output 3 modes of paths, for both policies, we execute the path with the lowest predicted collision cost. Attitudes are set to face towards the goal locations.
The results are recorded in Table 1. Note that the path elevation we report is the planned $z$ coordinates of the path rather than the relative elevation with respect to the terrain. As a reference, the average $z$ coordinate of the terrain is $18.41$m. We observe that our policy is able to generate path plans with lower altitude than BC, while having similar path lengths. The average depth prediction error, after converting from log scale to the original scale (in meters), is 1.82m. The average inference time is $0.0123$s with a standard deviation of $0.0009$. Note that, out of the 20 pairs of start and goal locations, our method successfully produced collision-free paths except for one pair of start and goal locations where the view at the start location is highly occluded by the mountain in front of the camera.
@@
Visualization of a comparison between the path generated by the expert and the path generated by our policy is shown in Fig 1. Note that our policy is able to generate paths lower than the expert demonstrations, since we supervised the elevation of the path planes directly using the terrain.
@@ 4 Conclusions
In this paper, we proposed a method to train a policy that’s capable of low altitude path planning with only RGB images and attitude information. Our method combines BC and self-supervised learning to enable the learned policy to outperform student policy trained with standard BC approach.
One limitation of the current method is that it does not take into account the max climb rate of the aircraft during training. Potential solutions include adding a differentiable optimization-based tracking controller (e.g. differentiable MPC) after the student policy and train the composed modules end-to-end as done in [26].

### v2

@@ 4 Simulation Experiments
Figure 6: Example of collected images. At each planned vehicle pose, we collect two RGB images and a depth image. We collect a RGB image (left) and its corresponding depth image (right) from a forward facing camera, and another RGB image (middle) from a downward facing camera to capture extra information.
In this section, we explain the setup of the simulation experiments and present the results.
@@ 4.1 Terrain
We build a custom terrain using Blender [27]. The terrain with the corresponding point cloud are visualized in Fig 5, The terrain is exported to NVIDIA Isaac Sim to perform simulation experiments. The size of the terrain is about 820m $\times$ 820m $\times$ 60m, which is scaled down from its original size to accelerate the rendering and data collection process.
@@ 4.2 Expert Planner Parameters
We set $\alpha=1.0,\beta=0.1$ in Eq 1. The minimum elevation with respect to the terrain is set to $5.0$m and the maximum elevation is set to $15.0$m. Since our terrain is scaled down, we accordingly scale down the minimum turning radius to $3$m and set the maximum climbing rate to $0.6$ rad/s. The cruise speed is set to $5$m/s. The range for RRT* is decreased to $300$m. The resolution used to check state validity is increased to $0.01$.
@@ 4.3 Dataset Preparation
We collect expert demonstrations on 1/4 of the terrain and use the full terrain for testing. We collect 1200 trajectories, randomly sampling start and goal locations such that each pair of start and goal locations are separated by at least 50m.
@@
To keep the inputs at similar scale, the heading vectors are divided by $h_{\text{const}}=100.0$, the log depth images are divided by $10.0$ and RGB images are divided by $255$.
@@ 4.5 Baseline Policy
In addition, we also train a baseline policy for comparison study, which uses the standard behavior cloning (BC) approach. In other words, the baseline policy does not have the loss terms ${\mathcal{L}}_{\text{terrain}}$ and ${\mathcal{L}}_{\text{elevation}}$, and the loss term ${\mathcal{L}}_{\text{bc}}$ applies to all three coordinates $x,y,z$ instead of only $x,y$. The other setups are the same (i.e. dataset, architecture, and hyperparameters).
@@ 4.6 Results
To compare with the baseline policy, which we term as BC, we select 20 pairs of start and goal locations that are separated by no more than 100m (since the policies were trained on heading vectors with norms less than the path length equivalent to $H$ steps, which is around $112.5$m). The start and goal locations are selected only in the unseen part of the terrain. Since both BC and our policy output 3 modes of paths, for both policies, we execute the path with the lowest predicted collision cost. The yaw angles of the camera are computed so that the camera always faces the goal location, and the roll and pitch angles are set to 0.
The results are recorded in Table 1. Note that the path elevation we report is the planned $z$-coordinates of the path rather than the relative elevation with respect to the terrain. As a reference, the average $z$-coordinate of the terrain is $18.41$m. We observe that our policy is able to generate path plans with lower altitude than BC, while having similar path lengths. The average depth prediction error, after converting from log scale to the original scale (in meters), is 1.82m. The average inference time is $0.0123$s with a standard deviation of $0.0009$ on an Nvidia GeForce RTX 4090 GPU. Note that out of the 20 pairs of start and goal locations, our method successfully produced collision-free paths for all but one pair of start and goal locations, where the view at the start location is highly occluded by the mountain in front of the camera.
@@
Visualization of a comparison between the path generated by the expert and the path generated by our policy is shown in Fig 1. Note that our policy is able to generate paths lower than the expert demonstrations, since we supervised the elevation of the path plans directly using the terrain.
@@ 5 Conclusions
In this paper, we proposed a method to train a policy that is capable of low-altitude path planning with only RGB images and attitude information. Our method combines BC and self-supervised learning to enable the learned policy to outperform the student policy trained with only a standard BC approach.
One limitation of the current method is that it does not explicitly take into account the max climb rate of the aircraft during training. Potential solutions include adding a differentiable optimization-based tracking controller (e.g. differentiable MPC) after the student policy and training the composed modules end-to-end as done in [29].